OpenStudy (fanduekisses):

Consider the curve defines by 2y^3+6x^2y-12x^2+6y=1

OpenStudy (fanduekisses):

\[2y^{3}+6x^{2}y-12x^{2}+6y=1\]

OpenStudy (fanduekisses):

Show that \[\frac{ dy }{ dx }=\frac{ 4x-2xy }{ x^{2}+y^{2}+1 }\]

OpenStudy (anonymous):

Do you know how to implicitly differentiate?

OpenStudy (fanduekisses):

kinda, I put dy/dx when there is a y

OpenStudy (anonymous):

\[2 y^3+6 x^2y-12 x^2+6y=1\] diff. w.r.t. x \[6 y^2\frac{ dy }{ dx }+6\left( 2xy+x^2\frac{ dy }{ dx } \right)-24x+6\frac{ dy }{ dx }=0\] divide by 6 and then find dy/dx

OpenStudy (fanduekisses):

How did you get 6(2xy+x2dydx)

OpenStudy (fanduekisses):

Did you factor?

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left( uv \right)=u \frac{ dv }{ dx }+v \frac{ du }{ dx }\]

OpenStudy (fanduekisses):

You used the product rule then factored?

OpenStudy (anonymous):

keep 6 outside then differentiate \[x^2y\]

OpenStudy (fanduekisses):

So after dividing I get: So after dividing I get: \[6y^2+2xy+x^{2}\frac{ dy }{ dx }-24x+6\frac{ dy }{ dx }=0\]

OpenStudy (fanduekisses):

Subtract the 2xy now?

OpenStudy (fanduekisses):

@campbell_st

OpenStudy (fanduekisses):

Ok so I get \[x^2\frac{ dy }{ dx }+6\frac{ dy }{ dx }=-6y^2-2xy-24x\]

OpenStudy (anonymous):

\[y^2\frac{ dy }{ dx }+2 xy+x^2\frac{ dy }{ dx }-4x+\frac{ dy }{ dx }=0\] \[\left( y^2+x^2+1 \right)\frac{ dy }{ dx }=4x-2xy\] \[\frac{ dy }{ dx }=?\]

OpenStudy (fanduekisses):

Am I doing it right so far?

OpenStudy (fanduekisses):

Where did the dy/dx next to the y^2 come from?

OpenStudy (fanduekisses):

Ohh I see I forgot

OpenStudy (fanduekisses):

Then divide everythign inside the parentheses to both sides

OpenStudy (fanduekisses):

ohh I got it now hehe thanks ^_^

OpenStudy (fanduekisses):

Now it asks me to write an equation of each horizontal tangent line to the curve, what exactly does that mean? I'm confused

OpenStudy (anonymous):

put \[\frac{ dy }{ dx }=0\]